An aluminium can of cylindrical shape contains 500 cm^3 of water. The area of the inner cross section of the can is 125 cm^2. All measurements refer to 10°C. Find the rise in the water level if the temperature increase to 80°C. The coefficient of linear expansion of aluminium= 23x10^-6/°C and average coefficient of volume expansion of water = 3.2x10^-4/°C.


Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

increase in volume of water=(3.2*10^-4)*70, area expansion of aluminium=(2*23*10^-6)*70, rise in water level=

(new volume of water)/(new area of cross section)-4cm